A company makes an investment in a new delivery van costing $47,000 and is expected to have a service life of 200,000 miles at an annual usage of 40,00 miles per year. The delivery service is expect to generate the following profits: $15K, in odd years and, - $20K, in even years, and the project is evaluated at a rate of 6%, A) What is the present value of this investment? B) What is the annual equivalent?
Answer #1

Useful life of investment=5 years

Initial cost=47000

Profits in year 1,3 &5=$15K

Profits in year 2,4=$20K

Ans a)

Present worth of Investment=PW of Profits-Initial costs=(15000/1.06)+(20000/1.06^2)+(15000/1.06^3)+(20000/1.06^4)+(15000/1.06^5)-47000=71595.90-47000=24595.9

Ans 2)

Annual equivalent=Present worth/Sum of discounting factors=24595.9/4.2125=$5838.98